# Diophantine Graphs

### v1

@@ 4.1 The Degree of a Vertex in $D_{n}$
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u\in V(D_{n})$, $p$ is a prime number,
$\deg(u)=n-1\quad\mbox{if and only if}\quad f(u)\mid n\quad\mbox{{or}}\quad\frac{n}{2}<f(u)=p^{\acute{v}_{p}(n)}<n,$
@@
Assume that $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$ with $\deg(u)=n-1$. By the way of contradiction, suppose
$f(u)\nmid n\quad\mbox{and}\quad\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{or}\quad\frac{n}{2}\nless f(u)=p^{\acute{v}_{p}(n)}\nless n\right),$
where $p$ is a prime number. In case of $f(u)=p^{\acute{v}_{p}(n)}$ and $1<f(u)<\frac{n}{2}$, we have
$(f(u),f(v))=(f(u),2f(u))=\left(p^{\acute{v}_{p}(n)},2p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
where $v\in V(D_{n})$ such that $f(v)=2f(n)<n$, which contradicts the full degree of $u$. In the other case, when $f(u)\neq p^{\acute{v}_{p}(n)}$ and $f(u)\nmid n$, we have that there exists a prime number $p\mid f(u)$ such that $v_{p}(f(u))>v_{p}(n).$ Then we get
$v_{p}\left(f(u)\right)\geq v_{p}(n)+1=\acute{v}_{p}(n),$
so $p^{\acute{v}_{p}(n)}\mid f(u)$. Therefore, one can see that
$\left(f(u),p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
which is a contradiction of the full degree of $u$. Therefore, one can see that
$f(u)\nmid n\quad\mbox{{or}}\quad f(u)=p^{\acute{v}_{p}(n)},\quad\frac{n}{2}<p^{\acute{v}_{p}(n)}<n.$
Conversely, in the first case of $f(u)\mid n$, we have for each $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the second case of $f(u)=p^{\acute{v}_{p}(n)}$ and $\frac{n}{2}<f(u)<n$, where $p$ is a prime number, we get any proper multiple of $f(u)$ is bigger than $n$. Since $f(u)=p^{\acute{v}_{p}(n)}$, therefore for all $v\in V(D_{n})$, $(f(u),f(v))\mid p^{v_{p}(n)}$, which means for all $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the both cases, we get $\deg(u)=n-1$. ∎
Similar results can be deduced in the case of the maximal prime graphs $R_{n}$ for Lemma 3.1 and Theorem 4.1. The properties of the maximal Diophantine graphs $D_{n}$ given in Lemma 3.1 and Theorem 4.1 are essential in what follows.
@@
Let $D_{p^{k}}$ be the maximal Diophantine graphs with $p^{k}$ vertices, where $k>1$ and $p$ is a prime number. $D_{p^{k}}$ can not be a prime graph because the number full degree vertices in $R_{p^{k}}$ that is equal to
@@
is less than of the number full degree vertices in $D_{p^{k}}$ which is at least
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $f^{*}(u)=p^{k}$, for some $p$ is a prime number and $k\in\mathbb{N}$, then
@@
Let $D_{n}$ be the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u\in V(D_{n})$ such that $f^{*}(u)=p^{k}$, where $p$ is a prime number and $k\geq 0$. In the case of $k=0$, we have $f^{*}(u)=1$. Since $f^{*}(u)=\frac{f(u)}{(f(u),n)}$, we have $f(u)\mid n$. Therefore, using Theorem 4.1, $\deg(u)=n-1$. In the other case of $k>0$, we have
@@
Therefore,
$\deg(u)=n-1-\Big{(}\Big{|}M_{p^{\acute{v}_{p}(n)}}\Big{|}-1\Big{)}=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor.$
Hence, we have the proof. ∎
Example 4.1.
In $D_{20}$, let $f(u)=18$ be the label of a vertex $u\in D_{20}$. Then $f^{*}(u)=9=3^{2}$ and we have $k=2>0$. Therefore, Using Lemma 4.1, we get the following
$\deg(u)=20-\left\lfloor\frac{20}{3^{\acute{v}_{3}(20)}}\right\rfloor=14.$
Another way to compute the degree of $u$ with label $18$. The neighborhood of $u$ is the set of adjacent vertices of $u$ that has labels belong to the set $\{1,2,4,5,7,8,10,11,13,14,16,17,19,20\}$. Therefore, $\deg(u)=14$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph with order $n$, $f$ is a Diophantine labeling and $u\in V(D_{n})$. If $f^{*}(u)=\prod\limits_{i=1}^{r}p_{i}^{k_{i}}$, where $p_{i}$, $i=1,2,\dots,r$ are distinct prime numbers and $k_{i}\geq 0$, $i=1,2,\dots,r$, then
$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$f^{*}(u)=1.$}\\ n-\sum\limits_{1\leq i\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\bigg{\rfloor}+\sum\limits_{1\leq i.j\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\bigg{\rfloor}-\dots+(-1)^{r}\Bigg{\lfloor}\frac{n}{\prod\limits_{1\leq i\leq r}p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},&\hbox{$f^{*}(u)>1.$}\end{array}\right.$
Similar results of Theorems 3.1, 4.2 can be obtained in the maximal prime graph $R_{n}$, by putting $v_{p}(n)=0$.
4.2 Vertices of Equal Degrees $<n-1$
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$. If $f(u)\mid f(v)$, then $N(u)\supseteq N(v)$, where $N(s)$ defines the neighborhood of $s$ as the set of all vertices in $D_{n}$ that join the vertex $s$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$. If $\deg(u)=\deg(v)$, then $N(u)=N(v)$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that
$f(u)\mid f(v)\quad\mbox{and}\quad\deg(u)=\deg(v).$
Let $s\in V(D_{n})-\{u,v\}$ such that $sv\in E(D_{n})$. Then, using Lemma 4.2, $su\in E(D_{n})$. Hence, $N(u)\supseteq N(v)$. On the other hand, let $s\in V(D_{n})-\{u,v\}$ such that $su\in E(D_{n})$. Suppose by contrary that $sv\notin E(D_{n})$. Since $f(u)\mid f(v)$ and using Lemma 4.2, therefore $N(u)\supset N(v)$. Then we have $\deg(u)>\deg(v)$, which contradicts the hypothesis of $\deg(u)=\deg(v)$. Therefore, $sv\in E(D_{n})$. Consequently, $N(u)\subseteq N(v)$. Hence, $N(u)=N(v)$. ∎
It is clear that, if $N(u)=N(v)$, then $\deg(u)=\deg(v)$. Thus, under the assumption that $f(u)\mid f(v)$, we have
$N(u)=N(v)\quad\mbox{ if and only if}\quad\deg(u)=\deg(v).$
Furthermore, the following example shows that the condition $f(u)\mid f(v)$ is necessary for Lemma 4.3.
@@
In $D_{20}$, let $f(u)=7$ and $f(v)=8$, where $u,v\in V(D_{20})$. $N(u)=V(D_{20})-\{u_{1},u\}$, where $u_{1}\in V(D_{20})$ such that $f(u_{1})=14$ and $N(v)=V(D_{20})-\{v_{1},v\}$, where $v_{1}\in V(D_{20})$ such that $f(v_{1})=16$. Then we have $\deg(u)=\deg(v)$ while $f(u)\nmid f(v)$, but $N(u)\neq N(v)$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $deg(u)<n-1$, then there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u\in V(D_{n})$ with $deg(u)<n-1$. Then there exists a vertex $s\in V(D_{n})$ such that $su\notin E(D_{n})$. Therefore, using Lemma 3.1, there exists a prime number $p$ such that $f(s),f(u)\in M_{p^{\acute{v}_{p}(n)}}$. Hence, there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$. ∎
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\geq 0$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p,q$ are two prime numbers. Then we have
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k>0$, where $p,q$ are primes numbers and $u$ $v\in V(D_{n})$ with $deg(u)=deg(v)$, $deg(u)<n-1$, $deg(v)<n-1$. Therefore, using Lemma 4.4, we have
@@
Consequently, $\acute{v}_{p}(n)\leq t$ and $\acute{v}_{q}(n)\leq k$. By using Lemma 4.1, we get
@@
Hence, we obtain the following equation
$\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Conversely, let $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$ and $\acute{v}_{p}(n)\leq t$, $\acute{v}_{q}(n)\leq k$. Then we get $p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{and}\ q^{\acute{v}_{q}(n)}\mid f(v).$ Thus, using Lemma 4.1, we have the following two equations
@@
Hence, the result follows. ∎
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices with $\deg(u)<n-1$, $\deg(v)<n-1$ such that $f(u)=p^{t}$, $f(v)=p^{k}$ for some $t,k\geq 0$, where $p$ is a prime number. Then we have
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices such that $f(v)=tf(u)$ for some $t\geq 1$ and $f(v)$ is not a prime power. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if for every prime number}\ p\mid t,\quad\acute{v}_{p}(n)\leq v_{p}(f(u))\quad\mbox{{or}}\quad\acute{v}_{p}(n)>v_{p}((f(v))).$
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $f(v)$ is not a prime power. Proof by contrapositive, suppose that there exists a prime number $p_{0}\mid t$ such that
$\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)>v_{p_{0}}((f(v))),$
@@
The first case contradicts the assumptions $f(u)\mid f(v)$. Therefore, let us consider the second case, which is
@@
Let a vertex $s\in V(D_{n})$ such that $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Therfore, using equation (4), $p_{0}^{\acute{v}_{p_{0}}(n)}\nmid p_{0}^{\acute{v}_{p_{0}}(f(u))}$. Since $p_{0}^{\acute{v}_{p_{0}}(f(u))}\mid f(u)$, we get $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)$. Consequently, $s\neq u$. Also, $s\neq v$ for $f(v)$ is not a prime power. Moreover, using equation (4), we get
@@
Thus,
$su\in E(D_{n})\quad\mbox{while}\quad sv\notin E(D_{n}).$
Since $f(u)\mid f(v)$ and using Lemma 4.2, we have $N(u)\supset N(v)$. Hence, $\deg(u)\neq\deg(v)$.
Conversely, suppose that, by contrapositive, $deg(u)\neq deg(v)$. Then, using Lemma 4.2, $N(u)\supset N(v)$. So, there exists a vertex $s\in V(D_{n})-\{u,v\}$ such that
$s\in N(u)\quad\mbox{and}\quad s\notin N(v).$
Consequently,
@@
$f(v),f(s)\in M_{p_{0}^{\acute{v}_{p_{0}}(n)}}\quad\mbox{and}\quad p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u).$ (6)
Let $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation 6, we have $s\neq u$. Since $f(v)=tf(u)$, we get $p_{0}\mid t$ and $s\neq v$. Therefore, using equation 5, we have the following
@@
Hence, using equation 7, we get $v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)$ and $v_{p_{0}}(f(v))\geq\acute{v}_{p_{0}}(n)$ for some prime number $p_{0}\mid t$. ∎

### v2

@@ 4 Degrees of Vertices
In this section, we proved that there are infinitely many Diophantine graphs which are not prime graphs using a characterization of formula for full degree vertices. Moreover, we give a formula for the degree of vertices. Necessary and sufficient conditions of equal degree vertices with specific labels and specific reduced labels are investigated.
@@ 4.1 The Degree of a Vertex in $D_{n}$
For each vertex $u$ in $D_{n}$ and prime number $p$, $\deg(u)=n-1$ if and only if
$f(u)=p^{\acute{v}_{p}(n)},\ \frac{n}{2}<f(u)<n,\quad\mbox{{or}}\quad f(u)\mid n,$
@@
Let $u\in V(D_{n})$ with $\deg(u)=n-1$. By the way of contradiction, suppose
$f(u)=p^{\acute{v}_{p}(n)},\ \frac{n}{2}<f(u)<n,\quad\mbox{and}\quad f(u)\mid n,$
or
$\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{or}\quad 1\leq f(u)\leq\frac{n}{2}\right)\quad\mbox{and}\quad f(u)\nmid n,$
where $p$ is a prime number. The first case is clearly a contradictory statement. The second case has two subcases as follows.
$\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{and}\quad f(u)\nmid n\right)\quad\mbox{or}\quad\left(1<f(u)<\frac{n}{2}\quad\mbox{and}\quad f(u)\nmid n\right),$
In case of $f(u)\neq p^{\acute{v}_{p}(n)}$ and $f(u)\nmid n$, we get that there exists a prime number $p$ such that $p\mid f(u)$ and $v_{p}(f(u))>v_{p}(n).$ Then, we obtain $v_{p}\left(f(u)\right)\geq\acute{v}_{p}(n),$ so $p^{\acute{v}_{p}(n)}\mid f(u)$. Consequently, one can see that $\left(f(u),p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$ which is a contradiction of the full degree of $u$. In the other case of $1<f(u)=p^{\acute{v}_{p}(n)}<\frac{n}{2}$, we have $(f(u),2p^{\acute{v}_{p}(n)})\nmid n,$ which contradicts the full degree of $u$. Hence, one can see that
$f(u)=p^{\acute{v}_{p}(n)},\ \frac{n}{2}<p^{\acute{v}_{p}(n)}<n\quad\mbox{{or}}\quad f(u)\mid n.$
Conversely, in the first case of $f(u)\mid n$, we have that for each $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the second case of $f(u)=p^{\acute{v}_{p}(n)}$ and $\frac{n}{2}<f(u)<n$, where $p$ is a prime number, we get that any proper multiple of $f(u)$ is greater than $n$. Since $f(u)=p^{\acute{v}_{p}(n)}$, therefore for all $v\in V(D_{n})$, $(f(u),f(v))\mid p^{v_{p}(n)}$, which means for all $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In both cases, we get $\deg(u)=n-1$. ∎
Similar results can be deduced in the case of a maximal prime graphs $R_{n}$ for Lemma 3.1 and Theorem 4.1. In light of Lemma 3.1, Theorem 4.1, and the following theorem 4.2 due to M. A. Seoud and M. Z. Youssef, we have the following results.
Theorem 4.2.
[11], [13] If $G$ is a simple graph of $n$ vertices which has more than
$\pi\left(n\right)+\pi\left(\frac{n}{2}\right)+1$
vertices of degree n1n-1, then $G$ is not prime, where $\pi(n)$ is the number of prime numbers not exceeding $n$.
@@
Let $p$ be a prime number and an integer $k>1$. Maximal Diophantine graphs $D_{p^{k}}$ with $p^{k}$ vertices can not be prime graphs because the number of full degree vertices in $R_{p^{k}}$ that is equal to
@@
and it is less than the number of full degree vertices in $D_{p^{k}}$ which is at least
@@
Let $u$ be a vertex in $D_{n}$. If $f^{*}(u)=p^{k}$ for some prime number $p$ and $k\in\mathbb{N}$, then
@@
Let $f^{*}(u)=p^{k}$, where $u\in V(D_{n})$, $p$ is a prime number and $k\in\mathbb{N}$. In the case of $k=0$, we have $f^{*}(u)=1$. Since $f^{*}(u)=\frac{f(u)}{(f(u),n)}$, we have $f(u)\mid n$. Therefore, using Theorem 4.1, $\deg(u)=n-1$. In the other case of $k>0$, we have
@@
Therefore, $\deg(u)=n-1-\left(\left|M_{p^{\acute{v}_{p}(n)}}\right|-1\right)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor.$
@@
Let $u$ be a vertex in $D_{n}$. If $f^{*}(u)=\prod\limits_{i=1}^{r}p_{i}^{k_{i}}$, where $p_{i}$, $i=1,2,\dots,r$ are distinct prime numbers and $k_{i}\in\mathbb{N}$, $i=1,2,\dots,r$, then
$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\\ \hskip 284.52756pt\hbox{if $f^{*}(u)=1.$}\\ n-\sum\limits_{1\leq i\leq r}\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor+\sum\limits_{1\leq i.j\leq r}\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,&\\ \hskip 284.52756pt\hbox{if $f^{*}(u)>1.$}\end{array}\right$.
Results similar to Theorems 3.1, 4.3 can be obtained in a maximal prime graph $R_{n}$, by putting $v_{p}(n)=0$.
4.2 Vertices of Equal Degrees <n1<n-1
@@
Let $u,v$ be two vertices in $D_{n}$. If $f(u)\mid f(v)$, then $N(u)-\{v\}\supseteq N(v)-\{u\}$.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)$. If $\deg(u)=\deg(v)$, then $N(u)-\{v\}=N(v)-\{u\}$.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)\ \mbox{and}\ \deg(u)=\deg(v).$ Let $s\in V(D_{n})-\{u,v\}$ such that $sv\in E(D_{n})$. Then, using Lemma 4.2, $su\in E(D_{n})$. Hence, $N(u)-\{v\}\supseteq N(v)-\{u\}$. On the other hand, let $s\in V(D_{n})-\{u,v\}$ such that $su\in E(D_{n})$. Suppose by contrary that $sv\notin E(D_{n})$. Since $f(u)\mid f(v)$ and using Lemma 4.2, therefore $N(u)-\{v\}\supset N(v)-\{u\}$. Then we have $\deg(u)>\deg(v)$, which contradicts the assumption of $\deg(u)=\deg(v)$. Therefore, $sv\in E(D_{n})$. Consequently, $N(u)-\{v\}\subseteq N(v)-\{u\}$. Hence, $N(u)-\{v\}=N(v)-\{u\}$. ∎
It is clear that if $N(u)-\{v\}=N(v)-\{u\}$, then $\deg(u)=\deg(v)$. Thus, under the assumption that $f(u)\mid f(v)$, we have $N(u)-\{v\}=N(v)-\{u\}$ if and only if $\deg(u)=\deg(v).$ Furthermore, the following example shows that the condition $f(u)\mid f(v)$ is necessary for Lemma 4.3.
@@
In $D_{20}$, let $f(u)=7$ and $f(v)=8$, where $u,v\in V(D_{20})$. Then $N(u)=V(D_{20})-\{u_{1},u\}$, where $u_{1}\in V(D_{20})$ such that $f(u_{1})=14$ and $N(v)=V(D_{20})-\{v_{1},v\}$, where $v_{1}\in V(D_{20})$ such that $f(v_{1})=16$. Then we have $\deg(u)=\deg(v)$, while $f(u)\nmid f(v)$, but $N(u)-\{v\}\neq N(v)-\{u\}$.
@@
Let $u$ be a vertex in $D_{n}$. If deg(u)<n1deg(u)<n-1, then there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$.
@@
Let $u$ be a vertex in $D_{n}$ with deg(u)<n1deg(u)<n-1. Then there exists a vertex $s\in V(D_{n})$ such that $su\notin E(D_{n})$. Therefore, using Lemma 3.1, there exists a prime number $p$ such that $f(s),f(u)\in M_{p^{\acute{v}_{p}(n)}}$. Hence, there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$. ∎
@@
Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\in\mathbb{Z^{+}}$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p,q$ are prime numbers. Then we have
@@
Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\in\mathbb{Z^{+}}$, where $p,q$ are prime numbers and $u,v\in V(D_{n})$ with $deg(u)=deg(v)$, deg(u)<n1deg(u)<n-1 and deg(v)<n1deg(v)<n-1. Therefore, using Lemma 4.4, we have
@@
Consequently, $\acute{v}_{p}(n)\leq t$ and $\acute{v}_{q}(n)\leq k$. Using Lemma 4.1, we get
@@
Hence, we obtain the following equation $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Conversely, let $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$ and $\acute{v}_{p}(n)\leq t$, $\acute{v}_{q}(n)\leq k$. Then we get
$p^{\acute{v}_{p}(n)}\mid f(u)\quad\mbox{and}\quad q^{\acute{v}_{q}(n)}\mid f(v).$
Thus, using Lemma 4.1, we have the following two equations
@@
Hence, the proof follows. ∎
@@
Let $f(u)=p^{t}$ and $f(v)=p^{k}$ for some $t,k\in\mathbb{Z^{+}}$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p$ is a prime number. Then we have
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$ and $f(v)$ is not a prime power. Then we have $\deg(u)=\deg(v)$ if and only if for every prime number $p\mid t$,
$\acute{v}_{p}(n)\leq v_{p}(f(u))\quad\mbox{{or}}\quad\acute{v}_{p}(n)>v_{p}(f(v)).$
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$ and $f(v)$ is not a prime power. Proof by contraposition, suppose that there exists a prime number $p_{0}\mid t$ such that
$\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)>v_{p_{0}}(f(v)),$
@@
The first case contradicts assumptions $f(u)\mid f(v)$. Therefore, let us consider the second case, which is
@@
Let a vertex $s\in V(D_{n})$ such that $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation (4), we have $p_{0}^{\acute{v}_{p_{0}}(n)}\nmid p_{0}^{\acute{v}_{p_{0}}(f(u))}.$ Since $p_{0}^{\acute{v}_{p_{0}}(f(u))}\mid f(u)$, we get $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)$. Consequently, $s\neq u$. In addition, $s\neq v$, since $f(v)$ is not a prime power. Moreover, using equation (4), we get
@@
Thus, $su\in E(D_{n})\ \mbox{while}\ sv\notin E(D_{n}).$ Since $f(u)\mid f(v)$ and using Lemma 4.2, we have $N(u)-\{v\}\supset N(v)-\{u\}$. Hence, $\deg(u)\neq\deg(v)$.
Conversely, suppose that, by contrapositive, $deg(u)\neq deg(v)$. Then, using Lemma 4.2, $N(u)-\{v\}\supset N(v)-\{u\}$. So, there exists a vertex $s\in V(D_{n})-\{u,v\}$ such that $s\in N(u)-\{v\}\ \mbox{and}\ s\notin N(v)-\{u\}.$ Consequently,
@@
$p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)\quad\mbox{and}\quad f(v),f(s)\in M_{p_{0}^{\acute{v}_{p_{0}}(n)}}.$ (6)
Let $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation (6), we have $s\neq u$. Since $f(v)=tf(u)$, we get $p_{0}\mid t$ and $s\neq v$. Therefore, using equation (5), we have
@@
Hence, using equation (7), we get
$v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)\quad\mbox{and}\quad v_{p_{0}}(f(v))\geq\acute{v}_{p_{0}}(n)$
for some prime number $p_{0}\mid t$. ∎